# Hyperbolic Dataset Distillation

### v1

@@ Abstract
To address the computational and storage challenges posed by large-scale datasets in deep learning, dataset distillation has been proposed to synthesize a compact dataset that replaces the original while maintaining comparable model performance. Unlike optimization-based approaches that require costly bi-level optimization, distribution matching (DM) methods improve efficiency by aligning the distributions of synthetic and original data, thereby eliminating nested optimization. DM achieves high computational efficiency and has emerged as a promising solution. However, existing DM methods, constrained to Euclidean space, treat data as independent and identically distributed points, overlooking complex geometric and hierarchical relationships. To overcome this limitation, we propose a novel hyperbolic dataset distillation method, termed HDD. Hyperbolic space, characterized by negative curvature and exponential volume growth with distance, naturally models hierarchical and tree-like structures. HDD embeds features extracted by a shallow network into the Lorentz hyperbolic space, where the discrepancy between synthetic and original data is measured by the hyperbolic (geodesic) distance between their centroids. By optimizing this distance, the hierarchical structure is explicitly integrated into the distillation process, guiding synthetic samples to gravitate towards the root-centric regions of the original data distribution while preserving their underlying geometric characteristics. Furthermore, we find that pruning in hyperbolic space requires only 20% of the distilled core set to retain model performance, while significantly improving training stability. Notably, HDD is seamlessly compatible with most existing DM methods, and extensive experiments on different datasets validate its effectiveness. To the best of our knowledge, this is the first work to incorporate the hyperbolic space into the dataset distillation process. The code will be released upon acceptance.
@@ 1 Introduction
Recently, deep neural networks (DNNs) have demonstrated outstanding performance across a wide range of tasks. However, the continuous performance improvement has led to increasingly large datasets, which in turn have escalated storage costs and computational demands, emerging as key bottlenecks in the further advancement of deep learning. To address this issue, dataset distillation (DD) has been proposed wang2018datasetdistillation . By condensing the information of the original dataset, DD synthesizes a significantly smaller artificial dataset while striving to achieve comparable model performance. Beyond this, DD has also been widely applied in various domains, such as neural architecture search DM ; n1 ; n2 ; n3 , continual learning c1 ; c2 , and privacy protection p1 ; p2 ; p4 ; li2022compressed .
@@ 2 Related Works
Dataset Distillation. Existing DD methods can be broadly categorized into three categories: gradient matching, trajectory matching, and distribution matching li2022awesome ; lei2023survey ; yu2023review ; liu2025survey . Gradient matching g1 ; dsa seeks to preserve critical information by minimizing the discrepancy between the gradients induced by synthetic and original samples during model training. Trajectory matching t1 ; t2 ; guo2024datm ; cui2023scaling ; li2023ddpp ; li2024iadd achieves fine-grained knowledge transfer by aligning the training trajectories of network parameters. Distribution matching DM ; DM1 ; DM2 improves the representational capacity of synthetic samples by aligning their statistical distributions with those of original data in feature or activation spaces. Recently, generative-based dataset distillation gu2024efficient ; d4m ; su2024diffusion ; li2024generative ; li2025generative ; li2025diffusion and decoupling optimization-based methods yin2023sre2l ; RDED ; GVBSM have been proposed, accelerating advancements in the field of dataset distillation. In this work, we introduce hyperbolic space into dataset distillation by leveraging its inherent negative curvature to impose the tree-like hierarchy of the original dataset onto synthetic data, thereby offering a novel perspective to address the fundamental challenges in dataset distillation.
@@ 3.1 Preliminaries
Problem Definition. Consider a large-scale original dataset =$\mathcal{R}=\left\{(r_{i}^{\text{real}},t_{i}^{\text{real}})\right\}_{i=1}^{|\mathcal{R}|}$, where $r_{i}^{\text{real}}$ represents the $i$-th sample instance from the original dataset, $t_{i}^{\text{real}}$ represents the corresponding label of the sample $r_{i}^{\text{real}}$ in the original dataset, and $|\mathcal{R}|$ is the total number of samples in the original dataset. The goal of dataset distillation is to construct a significantly smaller synthetic dataset $\mathcal{S}=\left\{(s_{j}^{\text{syn}},t_{j}^{\text{syn}})\right\}_{j=1}^{|\mathcal{S}|}$, where $s_{j}^{\text{syn}}$ represents the $j$-th synthetic sample instance, $t_{j}^{\text{syn}}$ represents the corresponding label of the synthetic sample $s_{j}^{\text{syn}}$, and $|\mathcal{S}|$ is the total number of samples in the synthetic dataset, with $|\mathcal{S}|\ll|\mathcal{R}|$. Such that a model trained on $\mathcal{S}$ (denoted $\theta_{\text{syn}}$) exhibits performance comparable to one trained on $\mathcal{R}$ (denoted $\theta_{\text{real}}$) when evaluated on previously unseen samples. Formally, let $P_{T}$ denote the true data distribution and $\ell$ a loss function (e.g., cross-entropy), then the optimal synthetic dataset is obtained by minimizing the discrepancy in performance between $\theta_{\text{syn}}$ and $\theta_{\text{real}}$ as follows:
$\mathcal{S}^{\star}=\arg\min_{E_{(x,t)}\sim P_{T}}\left\|\ell\left(\theta_{\text{syn}}(x),t\right)-\ell\left(\theta_{\text{real}}(x),t\right)\right\|,$ (1)
where $(x,t)$ represents a sample pair drawn from the true data distribution $P_{T}$, with $x$ denoting the data instance and $t$ its corresponding label.
@@
:=$\mathcal{L}:=\left\{\mathbf{x}\in\mathbb{R}^{n+1}\;\middle|\;\langle\mathbf{x},\mathbf{x}\rangle_{\mathcal{L}}=\frac{1}{K},\;x_{t}>0\right\}.$ (3)
@@ 3.2 Hyperbolic Dataset Distillation for Distribution Matching
Given the original dataset =$\mathcal{R}=\left\{(r_{i}^{\text{real}},t_{i}^{\text{real}})\right\}_{i=1}^{|\mathcal{R}|}$ and the synthetic dataset for update $\mathcal{S}=\left\{(s_{j}^{\text{syn}},t_{j}^{\text{syn}})\right\}_{j=1}^{|\mathcal{S}|}$ (where $|\mathcal{S}|\ll|\mathcal{R}|$), we first encode the data through a frozen pre-trained encoder $\phi$, generating corresponding feature vectors $v_{i}^{\rm real}$ and $v_{j}^{\rm syn}$ as follows:
@@
$z_{i}^{\rm real}$ $=\exp_{p_{0}}\bigl{(}v_{i}^{\rm real}\bigr{)}=\cosh\bigl{(}\|v_{i}^{\rm real}\|\bigr{)}\,p_{0}\;+\;\sinh\bigl{(}\|v_{i}^{\rm real}\|\bigr{)}\,\frac{v_{i}^{\rm real}}{\|v_{i}^{\rm real}\|},$ (6)
$z_{j}^{\rm syn}$ $=\exp_{p_{0}}\bigl{(}v_{j}^{\rm syn}\bigr{)}=\cosh\bigl{(}\|v_{j}^{\rm syn}\|\bigr{)}\,p_{0}\;+\;\sinh\bigl{(}\|v_{j}^{\rm syn}\|\bigr{)}\,\frac{v_{j}^{\rm syn}}{\|v_{j}^{\rm syn}\|}.$ (7)
@@
where $z$ denotes a point in the Lorentzian hyperbolic space $\mathbb{L}_{K}^{n}$ over which the Riemannian mean is optimized, and the Lorentzian hyperbolic distance $d_{L}$ on the upper-sheet hyperboloid model is
@@ 3.3 Loss Contribution of Samples at Different Levels
In hyperbolic space, samples embedded at lower levels tend to better represent category prototypes. When calculating the centroid, hyperbolic space can effectively assign different weights to relatively lower-level and higher-level samples, meaning their contributions to the centroid vary in influence. To gain explicit insight into how each sample influences the alignment between the original dataset $\mathcal{R}$ and the synthetic dataset $\mathcal{S}$ in hyperbolic space, we adopt a tangent space approximation centered at the origin $o\in\mathbb{L}^{n}_{K}$.Since the centroids of the sets ($\bar{z}^{\rm real}$ and $\bar{z}^{\rm syn}$) are near the origin, this approximation is reasonably effective. For $Z^{\rm real}$ and $Z^{\rm syn}$, respectively define the hyperbolic radius (distance to the origin) of each sample as:
@@
$u_{i}=\frac{r_{i}^{\text{real}}-\cosh r_{i}\,o}{\sinh r_{i}},\qquad v_{j}=\frac{s_{j}^{\text{syn}}-\cosh s_{j}\,o}{\sinh s_{j}}.$ (16)
@@ 4.1 Experimental Setup
Dataset. We evaluated HDD on several standard benchmark datasets, including Fashion-MNIST fashion , SVHN svhn , CIFAR-10 cifar , CIFAR-100 cifar , and the larger-scale TinyImageNet tiny . Additionally, for hybrid architecture experiments, we utilized the ImageWoof subset of ImageNet imagenet , which features higher resolution images. Please refer to Appendix E for detailed information about the used datasets.
@@
Implementation Details. Our hyperparameter settings follow the design of the DM DM , IDM DM2 , and Dance DM1 architectures. We adopt the differentiable siamese augmentation dsa enhancement method used in prior works. The synthetic dataset is learned using SGD. For DM with HDD, we train for 20,000 iterations, while for IDM with HDD and Dance with HDD, we train for 10,000 iterations. For all experiments, we set the batch size to 256. Additionally, for different experiments, we use distinct hyperbolic curvature $K$, gradient scaling factor $\lambda$, and synthetic image learning rate $r$, as detailed in Appendix G. All experiments are conducted on one RTX A6000 Ada GPU.
@@ 4.2 Main Results
Table 1 presents a comparative evaluation of our method against prior approaches on Fashion-MNIST fashion , SVHN svhn , CIFAR-10 cifar , and CIFAR-100 cifar . The results for TinyImageNet tiny are provided in Appendix I. IDM augmented with HDD, which exploits the hierarchical inductive bias of hyperbolic space, consistently outperforms the baseline IDM across all benchmarks. Under the IPC = 1 setting, IDM with HDD achieves classification accuracies of 78.5% on FashionMNIST (+1.1%), 67.8% on SVHN (+2.5%), 47.0% on CIFAR-10 (+1.8%), and 25.3% on CIFAR-100 (+3.2%), demonstrating its superiority in low-data regimes. With IPC = 10, the proposed method attains 61.3% accuracy on CIFAR-10, a 4.0% improvement over IDM. Under IPC = 50, it yields gains of 2.4%, 2.5%, and 2.4% on SVHN, CIFAR-10, and CIFAR-100, respectively. Furthermore, DM with HDD also exhibits notable enhancements relative to DM: on SVHN, accuracy increases by 3.1% (IPC = 1) and 2.3% (IPC = 10), and on CIFAR-10 (IPC = 1) by 2.3%. We present some of the distilled images in Appendix J.
@@
Random 51.4$\pm$3.8 73.8$\pm$0.7 82.5$\pm$0.7 14.6$\pm$1.6 35.1$\pm$4.1 70.9$\pm$0.9 14.4$\pm$2.0 26.0$\pm$1.2 43.4$\pm$1.0 4.2$\pm$0.3 14.6$\pm$0.5 30.0$\pm$0.4
Herding 67.0$\pm$1.9 71.1$\pm$0.7 71.9$\pm$0.8 20.9$\pm$1.3 50.5$\pm$3.3 72.6$\pm$0.8 21.5$\pm$1.2 31.6$\pm$0.7 40.4$\pm$0.6 8.4$\pm$0.3 17.3$\pm$0.3 33.7$\pm$0.5
K-Center 66.9$\pm$1.8 54.7$\pm$1.5 68.3$\pm$0.8 21.0$\pm$1.5 14.0$\pm$1.3 20.1$\pm$1.4 21.5$\pm$1.3 14.7$\pm$0.7 27.0$\pm$1.4 8.3$\pm$0.3 7.1$\pm$0.2 30.5$\pm$0.3
Forgetting - - - 12.1$\pm$5.6 16.8$\pm$1.2 27.2$\pm$1.5 13.5$\pm$1.5 23.3$\pm$1.0 23.3$\pm$1.1 4.5$\pm$0.3 15.1$\pm$0.2 30.5$\pm$0.4
DC g1 70.5$\pm$0.6 82.3$\pm$0.4 83.6$\pm$0.4 31.2$\pm$1.4 76.1$\pm$0.6 82.3$\pm$0.3 28.3$\pm$0.5 44.9$\pm$0.5 53.9$\pm$0.5 12.8$\pm$0.3 25.2$\pm$0.3 -
DSA dsa 70.6$\pm$0.6 84.6$\pm$0.3 88.7$\pm$0.3 27.5$\pm$1.4 79.2$\pm$0.5 84.4$\pm$0.4 28.8$\pm$0.7 52.1$\pm$0.5 60.6$\pm$0.5 13.9$\pm$0.3 32.3$\pm$0.3 42.8$\pm$0.4
CAFE point1 77.1$\pm$0.9 83.0$\pm$0.4 84.8$\pm$0.4 42.6$\pm$3.3 75.9$\pm$0.6 81.3$\pm$0.3 30.3$\pm$1.1 46.3$\pm$0.6 55.5$\pm$0.6 12.9$\pm$0.3 27.8$\pm$0.3 37.9$\pm$0.3
CAFE+DSA point1 73.7$\pm$0.7 83.0$\pm$0.3 88.2$\pm$0.3 42.9$\pm$3.0 77.9$\pm$0.6 82.3$\pm$0.4 31.6$\pm$0.8 50.9$\pm$0.5 62.3$\pm$0.4 14.0$\pm$0.3 31.5$\pm$0.2 42.9$\pm$0.2
DCC DCC - - - 34.3$\pm$1.6 76.2$\pm$0.8 83.3$\pm$0.2 34.0$\pm$0.7 54.4$\pm$0.5 64.2$\pm$0.4 14.6$\pm$0.3 33.5$\pm$0.3 39.4$\pm$0.4
G-VBSM GVBSM - - - - - - - 46.5$\pm$0.7 54.3$\pm$0.3 16.4$\pm$0.7 38.7$\pm$0.2 45.7$\pm$0.4
DataDAMpoint2 - - - - - - 32.0$\pm$1.2 54.2$\pm$0.8 67.0$\pm$0.4 14.5$\pm$0.5 34.8$\pm$0.5 49.4$\pm$0.3
DM DM 70.7$\pm$0.6 83.4$\pm$0.1 88.1$\pm$0.6 21.9$\pm$0.4 72.8$\pm$0.3 82.6$\pm$0.3 26.4$\pm$0.3 48.5$\pm$0.6 62.2$\pm$0.5 11.4$\pm$0.3 29.7$\pm$0.3 43.0$\pm$0.4
DM with HDD 72.1$\pm$0.2 84.0$\pm$0.1 88.8$\pm$0.4 25.0$\pm$0.2 75.1$\pm$0.2 83.0$\pm$0.3 28.7$\pm$0.2 50.3$\pm$0.3 63.2$\pm$0.4 13.3$\pm$0.2 30.1$\pm$0.1 43.8$\pm$0.2
IDM DM2 77.4$\pm$0.3 82.4$\pm$0.2 84.5$\pm$0.1 65.3$\pm$0.3 81.0$\pm$0.1 85.2$\pm$0.3 45.2$\pm$0.5 57.3$\pm$0.3 67.2$\pm$0.1 22.1$\pm$0.2 44.7$\pm$0.3 46.5$\pm$0.4
IDM with HDD 78.5$\pm$0.2 83.8$\pm$0.2 86.4$\pm$0.3 67.8$\pm$0.2 84.0$\pm$0.2 87.6$\pm$0.1 47.0$\pm$0.1 61.3$\pm$0.1 69.7$\pm$0.2 25.3$\pm$0.2 45.4$\pm$0.1 48.9$\pm$0.3
Whole Dataset 93.5$\pm$0.1 95.4$\pm$0.1 84.8$\pm$0.1 56.2$\pm$0.3

### v2

@@ Abstract
To address the computational and storage challenges posed by large-scale datasets in deep learning, dataset distillation has been proposed to synthesize a compact dataset that replaces the original while maintaining comparable model performance. Unlike optimization-based approaches that require costly bi-level optimization, distribution matching (DM) methods improve efficiency by aligning the distributions of synthetic and original data, thereby eliminating nested optimization. DM achieves high computational efficiency and has emerged as a promising solution. However, existing DM methods, constrained to Euclidean space, treat data as independent and identically distributed points, overlooking complex geometric and hierarchical relationships. To overcome this limitation, we propose a novel hyperbolic dataset distillation method, termed HDD. Hyperbolic space, characterized by negative curvature and exponential volume growth with distance, naturally models hierarchical and tree-like structures. HDD embeds features extracted by a shallow network into the Lorentz hyperbolic space, where the discrepancy between synthetic and original data is measured by the hyperbolic (geodesic) distance between their centroids. By optimizing this distance, the hierarchical structure is explicitly integrated into the distillation process, guiding synthetic samples to gravitate towards the root-centric regions of the original data distribution while preserving their underlying geometric characteristics. Furthermore, we find that pruning in hyperbolic space requires only 20% of the distilled core set to retain model performance, while significantly improving training stability. Notably, HDD is seamlessly compatible with most existing DM methods, and extensive experiments on different datasets validate its effectiveness. To the best of our knowledge, this is the first work to incorporate the hyperbolic space into the dataset distillation process. The code is available at https://github.com/Guang000/HDD.
@@ 1 Introduction
Recently, deep neural networks (DNNs) have demonstrated outstanding performance across a wide range of tasks. However, the continuous performance improvement has led to increasingly large datasets, which in turn have escalated storage costs and computational demands, emerging as key bottlenecks in the further advancement of deep learning. To address this issue, dataset distillation (DD) has been proposed wang2018datasetdistillation . By condensing the information of the original dataset, DD synthesizes a significantly smaller artificial dataset while striving to achieve comparable model performance. Beyond this, DD has also been widely applied in various domains, such as neural architecture search DM ; t2 ; n2 ; n3 , continual learning c1 ; c2 , and privacy protection p1 ; p2 ; p4 ; li2022compressed .
@@ 2 Related Works
Dataset Distillation. Existing DD methods can be broadly categorized into three categories: gradient matching, trajectory matching, and distribution matching li2022awesome ; lei2023survey ; yu2023review ; liu2025survey . Gradient matching g1 ; dsa seeks to preserve critical information by minimizing the discrepancy between the gradients induced by synthetic and original samples during model training. Trajectory matching t1 ; t2 ; guo2024datm ; cui2023scaling ; li2023ddpp ; li2024iadd achieves fine-grained knowledge transfer by aligning the training trajectories of network parameters. Distribution matching DM ; DM1 ; DM2 improves the representational capacity of synthetic samples by aligning their statistical distributions with those of original data in feature or activation spaces. Recently, generative-based dataset distillation gu2024efficient ; d4m ; su2024diffusion ; li2024generative ; li2025generative ; li2025diffusion ; li2025diff ; ye2025igds and decoupling optimization-based methods yin2023sre2l ; RDED ; GVBSM have been proposed, accelerating advancements in the field of dataset distillation. In this work, we introduce hyperbolic space into dataset distillation by leveraging its inherent negative curvature to impose the tree-like hierarchy of the original dataset onto synthetic data, thereby offering a novel perspective to address the fundamental challenges in dataset distillation.
@@ 3.1 Preliminaries
Problem Definition. Consider a large-scale original dataset $\mathcal{R}=\left\{(r_{i}^{\text{real}},t_{i}^{\text{real}})\right\}_{i=1}^{|\mathcal{R}|}$, where $r_{i}^{\text{real}}$ represents the $i$-th sample instance from the original dataset, $t_{i}^{\text{real}}$ represents the corresponding label of the sample $r_{i}^{\text{real}}$ in the original dataset, and $|\mathcal{R}|$ is the total number of samples in the original dataset. The goal of dataset distillation is to construct a significantly smaller synthetic dataset $\mathcal{S}=\left\{(s_{j}^{\text{syn}},t_{j}^{\text{syn}})\right\}_{j=1}^{|\mathcal{S}|}$, where $s_{j}^{\text{syn}}$ represents the $j$-th synthetic sample instance, $t_{j}^{\text{syn}}$ represents the corresponding label of the synthetic sample $s_{j}^{\text{syn}}$, and $|\mathcal{S}|$ is the total number of samples in the synthetic dataset, with $|\mathcal{S}|\ll|\mathcal{R}|$. Such that a model trained on $\mathcal{S}$ (denoted $\theta_{\text{syn}}$) exhibits performance comparable to one trained on $\mathcal{R}$ (denoted $\theta_{\text{real}}$) when evaluated on previously unseen samples. Formally, let $P_{T}$ denote the true data distribution and $\ell$ a loss function (e.g., cross-entropy), then the optimal synthetic dataset is obtained by minimizing the discrepancy in performance between $\theta_{\text{syn}}$ and $\theta_{\text{real}}$ as follows:
$\mathcal{S}^{\star}=\arg\min_{E_{(p,t)}\sim P_{T}}\left\|\ell\left(\theta_{\text{syn}}(p),t\right)-\ell\left(\theta_{\text{real}}(p),t\right)\right\|,$ (1)
where $(p,t)$ represents a sample pair drawn from the true data distribution $P_{T}$, with $p$ denoting the data instance and $t$ its corresponding label.
@@
$\mathcal{L}:=\left\{\mathbf{x}\in\mathbb{R}^{n+1}\;\middle|\;\langle\mathbf{x},\mathbf{x}\rangle_{\mathcal{L}}=\frac{1}{K},\;x_{t}>0\right\}.$ (3)
@@ 3.2 Hyperbolic Dataset Distillation for Distribution Matching
Given the original dataset $\mathcal{R}=\left\{(r_{i}^{\text{real}},t_{i}^{\text{real}})\right\}_{i=1}^{|\mathcal{R}|}$ and the synthetic dataset for update $\mathcal{S}=\left\{(s_{j}^{\text{syn}},t_{j}^{\text{syn}})\right\}_{j=1}^{|\mathcal{S}|}$ (where $|\mathcal{S}|\ll|\mathcal{R}|$), we first encode the data through a frozen pre-trained encoder $\phi$, generating corresponding feature vectors $v_{i}^{\rm real}$ and $v_{j}^{\rm syn}$ as follows:
@@
$z_{i}^{\rm real}$ $=\exp_{p_{0}}\bigl(v_{i}^{\rm real}\bigr)=\cosh\bigl(\sqrt{-K}\|v_{i}^{\rm real}\|\bigr)\,p_{0}\;+\;\sinh\bigl(\sqrt{-K}\|v_{i}^{\rm real}\|\bigr)\,\frac{v_{i}^{\rm real}}{\sqrt{-K}\|v_{i}^{\rm real}\|},$ (6)
$z_{j}^{\rm syn}$ $=\exp_{p_{0}}\bigl(v_{j}^{\rm syn}\bigr)=\cosh\bigl(\sqrt{-K}\|v_{j}^{\rm syn}\|\bigr)\,p_{0}\;+\;\sinh\bigl(\sqrt{-K}\|v_{j}^{\rm syn}\|\bigr)\,\frac{v_{j}^{\rm syn}}{\sqrt{-K}\|v_{j}^{\rm syn}\|}.$ (7)
@@
where $z$ denotes a point in the Lorentzian hyperbolic space $\mathbb{L}_{K}^{n}$ over which the Riemannian mean is optimized, and the Lorentzian hyperbolic distance $d_{L}$ on the upper‐sheet hyperboloid model is
@@ 3.3 Loss Contribution of Samples at Different Levels
In hyperbolic space, samples embedded at lower levels tend to better represent category prototypes. When calculating the centroid, hyperbolic space can effectively assign different weights to relatively lower-level and higher-level samples, meaning their contributions to the centroid vary in influence. To gain explicit insight into how each sample influences the alignment between the original dataset $\mathcal{R}$ and the synthetic dataset $\mathcal{S}$ in hyperbolic space, we adopt a tangent space approximation centered at the origin $o\in\mathbb{L}^{n}_{K}$.Since the centroids of the sets ($\bar{z}^{\rm real}$ and $\bar{z}^{\rm syn}$) are near the origin, this approximation is reasonably effective. For $Z^{\rm real}$ and $Z^{\rm syn}$, respectively, define the hyperbolic radius (distance to the origin) of each sample as:
@@
$u_{i}={r_{i}^{\text{real}}-\cosh r_{i}\,o},\qquad v_{j}={s_{j}^{\text{syn}}-\cosh s_{j}\,o}.$ (16)
@@ 4.1 Experimental Setup
Dataset. We evaluated HDD on several standard benchmark datasets, including Fashion-MNIST fashion , SVHN svhn , CIFAR-10 cifar , CIFAR-100 cifar , and the larger-scale TinyImageNet tiny . Additionally, for hybrid architecture experiments, we utilized the ImageWoof subset of ImageNet imagenet , which features higher resolution images. Please refer to Appendix E for detailed information about the datasets used.
@@
Implementation Details. Our hyperparameter settings follow the design of the DM DM , IDM DM2 , and Dance DM1 architectures. We adopt the differentiable siamese augmentation dsa enhancement method used in prior works. The synthetic dataset is learned using SGD. For DM with HDD, we train for 20,000 iterations, while for IDM with HDD and Dance with HDD, we train for 10,000 iterations. For all experiments, we set the batch size to 256. Additionally, for different experiments, we use distinct hyperbolic curvature $K$, gradient scaling factor $\lambda$, and synthetic image learning rate $r$, as detailed in Appendix G. All experiments are conducted on one RTX A6000 Ada GPU, except for Section 4.5.
@@ 4.2 Main Results
Table 1 presents a comparative evaluation of our method against prior approaches on Fashion-MNIST fashion , SVHN svhn , CIFAR-10 cifar , and CIFAR-100 cifar . The results for TinyImageNet tiny are provided in Appendix I. IDM augmented with HDD, which exploits the hierarchical inductive bias of hyperbolic space, consistently outperforms the baseline IDM across all benchmarks. Under the IPC = 1 setting, IDM with HDD achieves classification accuracies of 78.5% on FashionMNIST (+1.1%), 67.8% on SVHN (+2.5%), 47.0% on CIFAR-10 (+1.8%), and 25.3% on CIFAR-100 (+3.2%), demonstrating its superiority in low-data regimes. With IPC = 10, the proposed method attains 61.3% accuracy on CIFAR-10, a 4.0% improvement over IDM. Under IPC = 50, it yields gains of 2.4%, 2.5%, and 2.4% on SVHN, CIFAR-10, and CIFAR-100, respectively. Furthermore, DM with HDD also exhibits notable enhancements relative to DM: on SVHN, accuracy increases by 3.1% (IPC = 1) and 2.3% (IPC = 10), and on CIFAR-10 (IPC = 1) by 2.3%. We present some of the distilled images in Appendix K.
@@
[…]